OpenStudy (anonymous):

PLEASE HELP!!! Use two unit multipliers to convert 6ft^2 to square inches.

OpenStudy (anonymous):

Sorry I'm not helping with this one. I'm not sure what your teacher means by unit multipliers in this question. I can convert it, but it might not include what they want.

OpenStudy (anonymous):

Unit multipliers are ways to turn a unit value (6ft^2) into a pure number...well sort of. This page explains it better: http://donnayoung.org/math/unit_multipliers.htm

OpenStudy (anonymous):

can u help me convert it?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

So one feet has 12 inches in it. \[\frac{1 feet}{12 inches}\] But you don't have one feet now do you? No you have 6 feet square (and two feets attached your leg) and the only way to remove that pesky feet is to divide it away using this unit multipler: \[\frac{12 inches}{1 ft}\] Let me show you how: \[6 ft^2 * \frac{12 inches}{1 ft} = 72 ft*inches\] Notice how those feets canceled themselves out? But that one unit multipler wasn't enough, so lets do it again. \[72 ft * \frac{12 inches}{1 ft} = 864 inches\] And wolframAlpha agrees with me: http://www.wolframalpha.com/input/?i=6+ft^2+in+inches^2

OpenStudy (anonymous):

Thx!